The difference of fourteen and three times a number is the same as the sum of the number and negative ten. Find the number.
Answer by checkley75(3666)

14-3X=X-10
-3X-X=-10-14
-4X=-24
X=-24/-4
X=6 ANSWER
PROOF
14-3*6=6-10
14-18=-4
-4=-4